OpenStudy (bekkah323):

when someone buys a ticket for an airline flight, there is a .0995 probability that, that person will not show up for their flight. Because of this the airlines will usually overbook for a flight to make sure all the seats are full. An agent for Air America wants to book 24 people on an airline that can only seat 22. If 24 people are booked find the probability that not enough seats will be available, or in other words, that 23 or more people would show up.

OpenStudy (agent0smith):

if X is the number of people who show up\[\Large P(X \ge 23) = P(X=23) + P(X=24)\] Now use the binomial formula to find P(X=23) and P(X=24)

OpenStudy (bekkah323):

so nCx*px*(1-p)n-x Would the n=24, is 24 my trial

OpenStudy (bekkah323):

what is the 0.995? Is that my probability so (0.995)23=0.891

OpenStudy (agent0smith):

0.0995 is the p

OpenStudy (agent0smith):

n=24 And you have to do it once for x=23 and x=24